the area of a square and rectangle are equal if the side of the square is 30 centimetre and the breadth of the rectangle is 20 cm find the length of...

Grade:11

1 Answers

Vikas TU
3 years ago
Dear student 
Area of square and rectangle are equal 
So S^2 = l*b 
30*30 = l*20 
l = 900/20
l = 45 centimetre 
Hope this will help